The ratio of boys to girls in a class is 3:1. There are 36 students in the class. How many more boys than girls are there?

Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52781)

.
x girls;

3x boys.


x + 3x = 36


4x = 36


x = 36/4 = 9.     It is the number of girls.


3x = 3*9 = 27.    It is the number of boys.


27 - 9 = 18.      It is your answer.

Solved.


Answer by Boreal(15235)

3:1 is ratio
3x:1x is same ratio
total number is 4x and 4x=36 x=9
27:9 or 18 more boys than girls.